# Cylindrical Gravitational Waves in Einstein-Aether Theory

### 30 Dec 2023 v1

@@ 8 Energy and Momentum of Cylindrical Gravitational Waves
The Einstein pseudotensor and the von Freud superpotential [43] obey the following relations.
@@
As said before, when evaluating the energy-momentum pseudotensor, the metric must be written in a coordinate system where the connection coefficients vanish like $\mathcal{O}(1/r)$ or faster in the asymptotic limit. Thus, we can see from equations (82)-(103) that the energy and momentum densities are finite and reasonable at infinity. From equations (42) and (46), we can notice that the quantities $H_{t}$ and $\psi_{t}$ are oscillatory functions of the time. Since the quantities $H_{r}$ and $\psi_{r}$ are also oscillatory in the radial coordinate and decreasing at infinity. Thus, we have finite values of the energy-momentum pseudotensors at infinity. We can notice that EA energy-momentum pseudotensors are completely different from the GR ones by construction, i.e., if we assume $c_{1}=c_{2}=c_{3}=c_{4}=$ all the EA energy-momentum pseudotensor components are zero.

### 26 Feb 2024 v2

@@ 8 Energy and Momentum of Cylindrical Gravitational Waves
The Einstein pseudotensor and the von Freud superpotential [44] obey the following relations
@@
As said before, when evaluating the energy-momentum pseudotensor, the metric must be written in a coordinate system where the connection coefficients vanish like $\mathcal{O}(1/r)$ or faster in the asymptotic limit. Thus, we can see from equations (80)-(101) that the energy and momentum densities are finite and reasonable at infinity. From equations (42) and (44), we can notice that the quantities $H_{t}$ and $\psi_{t}$ are oscillatory functions of the time. Since the quantities $H_{r}$ and $\psi_{r}$ are also oscillatory in the radial coordinate and decreasing at infinity. Thus, we have finite values of the energy-momentum pseudotensors at infinity.
We can notice that EA energy-momentum pseudotensors are completely different from the GR ones by construction, i.e., if we assume $c_{1}=c_{2}=c_{3}=c_{4}=$ all the EA energy-momentum pseudotensor components are zero.